# Low-Acceleration Gravitational Anomaly from Bayesian 3D Modeling of Wide Binary Orbits: Methodology and Results with Gaia DR3

### v1

@@ 4.3 Future direction of wide binary gravity research
stacked velocity profile $v_{p}$ $s$ pure binaries desired H12, C24a
@@
$\tilde{v}$-distribution histogram of $\tilde{v}$ $s$ (or $s/r_{\rm{M}}$) $s/r_{\rm{M}}$ desired, but $s$ used by many B18, P18, C24c
@@
Note. (1) Variables: $s=$ sky-projected separation, $r_{\rm{M}}=$ MOND radius (Equation (30)), $v_{p}=$ magnitude of the sky-projected relative velocity, $v_{c}(s)(\equiv\sqrt{G_{\rm{N}}M_{\rm{tot}}/{s}})=$ Newtonian circular velocity at $s$, $v(\equiv|\mathbf{v}|)=$ magnitude of the 3D relative velocity $\mathbf{v}$, $r=$ physical separation in 3D space. (2) References: H12 - Hernandez et al. (2012), C24a - Chae (2023a), C24c - Chae (2024c), B18 - Banik & Zhao (2018), P18 - Pittordis & Sutherland (2018), C23a - Chae (2023a), C23b - Chae (2023b) (Python codes)
@@ 5 Conclusion and follow-up work
    With 125 wide binaries in a high acceleration range (“Newtonian regime”) $0.080.02<s/r_{\rm{M}}<$ 0.08 ($10^{-7.9}\lesssim g_{\rm{N}}\lesssim 10^{-6.7}$ m s-2), the Bayesian inference returns $\Gamma=0.000\pm 0.011$. Thus, internal dynamics of binaries is consistent with Newtonian gravity in acceleration regimes $\gtrsim 100a_{0}$ as far as the first-tier orbital motions are concerned. No first-tier gravitational anomaly is observed in this regime.
@@
    With 76 wide binaries in an intermediate acceleration range (“transition regime”) $0.500.15<s/r_{\rm{M}}<$ 0.50 ($10^{-9.5}\lesssim g_{\rm{N}}\lesssim 10^{-8.5}$ m s-2), the Bayesian inference returns $\Gamma=0.057\pm 0.017$. Newtonian gravity is in tension with these wide binaries at a significance of $3.4\sigma$.
@@
    With 35 wide binaries in a low acceleration range (“MOND regime”) $2.860.50<s/r_{\rm{M}}<$ 2.86 ($10^{-11.0}\lesssim g_{\rm{N}}\lesssim 10^{-9.5}$ m s-2), the Bayesian inference returns $\Gamma=0.085\pm 0.040$. Newtonian gravity is in tension with these wide binaries at a significance of $2.1\sigma$. While the value of $\Gamma$ appears to be higher than in the transition regime, the statistical uncertainty is much larger.
@@
    With 111 wide binaries in a combined acceleration range (“transition+MOND regime”) $2.860.15<s/r_{\rm{M}}<$ 2.86 ($10^{-11.0}\lesssim g_{\rm{N}}\lesssim 10^{-8.5}$ m s-2), the Bayesian inference returns $\Gamma=0.063\pm 0.015$. Newtonian gravity is in tension with these wide binaries at a significance of $4.2\sigma$.
@@
    The inferred values of $\Gamma$ in the MOND and transition+MOND regimes are consistent with a generic value $\approx 0.07$ in the low acceleration ($\lesssim a_{0}$) regime predicted by classical theories of MOND gravity (Bekenstein & Milgrom, 1984; Milgrom, 2010). However, to test a specific theory accurately, numerical simulations of wide binary orbits in the theory are required.
@@
    The magnitude of the gravitational anomaly quantified by $\Gamma$ agrees well with the results from three statistical methods (Chae, 2023a, 2024a, 2024c; Hernandez et al., 2024a) based on much larger samples.
@@
The present results are limited by the relatively large uncertainties ($\sim 280$ m s-1) of $v_{r}$ that are $\approx 10$ times larger than the uncertainties of the sky-projected velocity components. At the time of this writing, for about 80 wide binaries in a low acceleration regime of $s/r_{\rm{M}}>0.4$, RVs of both stars are being measured with uncertainties $\lesssim 50$ m s-1. With these newly measured RVs, the precision of $\Gamma$ in the MOND regime will improve dramatically. In this study, stars brighter than $M_{G}=3.8$ were excluded to obtain a uniform sample of main-sequence stars by precluding possible contamination by bright subgiant or giant stars. However, it is interesting to consider stars brighter than $M_{G}=3.8$ in the future because brighter stars allow easier measurements of RVs and subgiant or giant stars can be easily excluded in the color-magnitude diagram.
@@ Acknowledgments
I have discussed this work with a number of excellent scientists. In particular, during the summer of 2024 I have visited three institutions and had interesting discussions of this work based on preliminary results. Those scientists include Stacy McGaugh and Tobias Mistele at Case Western Reserve University, Arthur Kosowsky, David Turnshek, and Andrew Zentner at the University of Pittsburgh, and Federico Lelli, Laura Magrini, and Luca Pasquini (a visitor there) at Arcetri Astrophysical Observatory. Luca Pasquini is further thanked for helpful conversations on radial velocities of stars. I thank Bill Janesh at CWRU for help with computing during the summer visit. I thank mathematicians Daniela Calvetti and Erkki Somersalo at CWRU for illuminating discussions on statistical consolidation. I thank Xavier Hernandez and Pavel Kroupa for discussions on wide binary gravity tests. I thank Byeong-Cheol Lee, Dongwook Lim, and David Turnshek for collaboration on measuring radial velocities of stars. I thank Yong Tian and Youngsub Yoon at Sejong for assistance and undertaking new statistical analyses of wide binaries. Finally, I thank Moti Milgrom for encouragements and inspirations of gravity research. Initial results of this work were presented by an invited talk given at the Korean Society of High Energy Physics 2024 Fall Meeting (https://indico.kshep.kr/event/839/contributions/4051). This work was supported by the National Research Foundation of Korea (grant No. NRF-2022R1A2C1092306).

### v3

@@ 4.3 Future direction of wide binary gravity research
stacked velocity profile $v_{p}$ $s$ pure binaries desired H12, H23, C24a, H25
@@
$\tilde{v}$-distribution histogram of $\tilde{v}$ $s$ (or $s/r_{\rm{M}}$) $s/r_{\rm{M}}$ desired, but $s$ used by many B18, P18, H24, C24c
@@
Note. (1) Variables: $s=$ sky-projected separation, $r_{\rm{M}}=$ MOND radius (Equation (31)), $v_{p}=$ magnitude of the sky-projected relative velocity, $v_{c}(s)(\equiv\sqrt{G_{\rm{N}}M_{\rm{tot}}/{s}})=$ Newtonian circular velocity at $s$, $v(\equiv|\mathbf{v}|)=$ magnitude of the 3D relative velocity $\mathbf{v}$, $r=$ physical separation in 3D space. (2) References: H12 - Hernandez et al. (2012), H23 - Hernandez (2023), C24a - Chae (2023a), H25 - Hernandez & Kroupa (2025), C24c - Chae (2024c), B18 - Banik & Zhao (2018), P18 - Pittordis & Sutherland (2018), H24 - Hernandez et al. (2024a), C23a - Chae (2023a), C23b - Chae (2023b) (Python codes)
@@ 5 Conclusion and follow-up work
    With 125 wide binaries in a high acceleration range (“Newtonian regime”) $0.080.02<s/r_{\rm{M}}<$ 0.08 ($10^{-7.9}\lesssim g_{\rm{N}}\lesssim 10^{-6.7}$ m s-2), the Bayesian inference returns $\Gamma=0.000\pm 0.011$ or $\gamma_{g}=1.00\pm 0.05$. Thus, internal dynamics of binaries is consistent with Newtonian gravity in acceleration regimes $\gtrsim 100a_{0}$ as far as the first-tier orbital motions are concerned. No first-tier gravitational anomaly is observed in this regime.
@@
    With 76 wide binaries in an intermediate acceleration range (“transition regime”) $0.500.15<s/r_{\rm{M}}<$ 0.50 ($10^{-9.5}\lesssim g_{\rm{N}}\lesssim 10^{-8.5}$ m s-2), the Bayesian inference returns $\Gamma=0.057\pm 0.017$ or $\gamma_{g}=1.30_{-0.09}^{+0.11}$. Newtonian gravity is in tension with these wide binaries at a significance of $3.4\sigma$.
@@
    With 35 wide binaries in a low acceleration range (“MOND regime”) $2.860.50<s/r_{\rm{M}}<$ 2.86 ($10^{-11.0}\lesssim g_{\rm{N}}\lesssim 10^{-9.5}$ m s-2), the Bayesian inference returns $\Gamma=0.085\pm 0.040$ or $\gamma_{g}=1.48_{-0.23}^{+0.30}$. Newtonian gravity is in tension with these wide binaries at a significance of $2.1\sigma$. While the value of $\Gamma$ appears to be higher than in the transition regime, the statistical uncertainty is much larger.
@@
    With 111 wide binaries in a combined acceleration range (“transition+MOND regime”) $2.860.15<s/r_{\rm{M}}<$ 2.86 ($10^{-11.0}\lesssim g_{\rm{N}}\lesssim 10^{-8.5}$ m s-2), the Bayesian inference returns $\Gamma=0.063\pm 0.015$ or $\gamma_{g}=1.34_{-0.08}^{+0.10}$. Newtonian gravity is in tension with these wide binaries at a significance of $4.2\sigma$.
@@
    The inferred values of $\Gamma$ in the MOND and transition+MOND regimes are consistent with a generic value $\Gamma\approx 0.07$ or $\gamma_{g}\approx 1.4$ in the low acceleration ($\lesssim a_{0}$) regime predicted by classical theories of MOND gravity (Bekenstein & Milgrom, 1984; Milgrom, 2010). However, to test a specific theory accurately, numerical simulations of wide binary orbits in the theory are required.
@@
    The magnitude of the gravitational anomaly quantified by $\Gamma$ agrees well with the results from three statistical methods (Chae, 2023a, 2024a, 2024c; Hernandez, 2023; Hernandez et al., 2024a; Hernandez & Kroupa, 2025) based on much larger samples.
@@
    Eight wide binaries with abnormal PDFs of $\Gamma$ are found (Table 2). If any of these does not have kinematic contaminants or abnormal RV data, it will individually rule out standard gravity. These systems require further studies.
  8. 8.
@@
The present results are limited by the relatively large uncertainties ($\sim 280$ m s-1) of $v_{r}$ that are $\approx 10$ times larger than the uncertainties of the sky-projected velocity components. At the time of this writing, for about 80 wide binaries in a low acceleration regime of $s/r_{\rm{M}}>0.4$, RVs of both stars are being measured with uncertainties $\lesssim 50$ m s-1. With these newly measured RVs, the precision of $\Gamma$ in the MOND regime will improve dramatically. In this study, stars brighter than $M_{G}=3.8$ were excluded to obtain a uniform sample of main-sequence stars by precluding possible contamination by bright subgiant or giant stars. However, as long as involved issues (such as higher probabilities of hiding close companions and availability of accurate masses) can be resolved, it is interesting to consider stars brighter than $M_{G}=3.8$ in the future because brighter stars allow easier measurements of RVs.
@@ Acknowledgments
I thank an anonymous referee for useful and constructive suggestions that led to improved and richer analyses and presentations. I have discussed this work with a number of excellent scientists. In particular, during the summer of 2024 I have visited three institutions and had interesting discussions of this work based on preliminary results. Those scientists include Stacy McGaugh and Tobias Mistele at Case Western Reserve University, Arthur Kosowsky, David Turnshek, and Andrew Zentner at the University of Pittsburgh, and Federico Lelli, Laura Magrini, and Luca Pasquini (a visitor there) at Arcetri Astrophysical Observatory. Luca Pasquini is further thanked for helpful conversations on radial velocities of stars. I thank Bill Janesh at CWRU for help with computing during the summer visit. I thank mathematicians Daniela Calvetti and Erkki Somersalo at CWRU for illuminating discussions on statistical consolidation. I thank Xavier Hernandez and Pavel Kroupa for discussions on wide binary gravity tests. I thank Byeong-Cheol Lee, Dongwook Lim, and David Turnshek for collaboration on measuring radial velocities of stars. I thank Yong Tian and Youngsub Yoon at Sejong for assistance and undertaking new statistical analyses of wide binaries. Y. Yoon is further thanked for carrying out independent calculations of the perspective effects based on the equations of Shaya & Olling (2011) with which the calculations of Appendix B agree well. Finally, I thank Moti Milgrom for encouragements and inspirations of gravity research. Initial results of this work were presented by an invited talk given at the Korean Society of High Energy Physics 2024 Fall Meeting (https://indico.kshep.kr/event/839/contributions/4051). This work was supported by the National Research Foundation of Korea (grant No. NRF-2022R1A2C1092306).
@@ Appendix A Relating the orbital plane to the observer’s frame with the Euler angles
Appendix B Perspective effects in proper motions and radial velocities
For a binary system, perspective effects (Shaya & Olling, 2011) refer to apparent relative motions between two stars, observed as biased PMs and RVs, due to the motion of the barycenter with respect to the Sun in the physical 3D space. Perspective effects are expected to be very small for wide binaries in the solar neighborhoods with projected separation $s<20$ kau (El-Badry, 2019). Nevertheless, here I describe a direct numerical algorithm and use it to investigate the perspective effects.
A binary is described using the notation introduced in Section 2.1. With R.A.’s and declinations of two stars $A$ and $B$, the barycenter’s azimuthal and polar angles $\theta$ and $\phi$ (which should not be confused with those used in Appendix A) in the equatorial frame are given by
$\phi=(\alpha_{A}+\alpha_{B})/2\text{\hskip 10.00002ptand\hskip 10.00002pt}\theta=\pi/2-(\delta_{A}+\delta_{B})/2,$ (B1)
where all angles are expressed in radians and I assume that the barycenter is the middle between the two stars for simplicity. The system velocity components in the R.A., delc., and radial directions are given by
$\left.\begin{array}[]{cl}V_{\alpha}=&4.7404\times 10^{-3}\times d_{M}(\mu^{\star}_{\alpha,A}+\mu^{\star}_{\alpha,B})/2\text{ km s}^{-1}\\ V_{\delta}=&4.7404\times 10^{-3}\times d_{M}(\mu_{\delta,A}+\mu_{\delta,B})/2\text{ km s}^{-1}\\ V_{r}=&({\rm{RV}}_{A}+{\rm{RV}}_{B})/2\text{ km s}^{-1}\end{array}\right\}.$ (B2)
In the geocentric Cartesian $XYZ$ frame the velocity components are
$\left.\begin{array}[]{cl}V_{X}=&V_{r}\sin\theta\cos\phi-V_{\delta}\cos\theta\cos\phi-V_{\alpha}\sin\phi\\ V_{Y}=&V_{r}\sin\theta\sin\phi-V_{\delta}\cos\theta\sin\phi+V_{\alpha}\cos\phi\\ V_{Z}=&V_{r}\cos\theta+V_{\delta}\sin\theta\end{array}\right\}.$ (B3)
Now I consider mock physical motions of two stars in the 3D space with the system velocity given by Equation (B3) for time $\Delta t$ into the future (or past). The current positions of the stars $A$ and $B$ are identified by subscript 0 while the future positions by 1. At the current time $t_{0}$, the azimuthal and polar coordinates are obtained by the relations $\phi=\alpha$ and $\theta=\pi/2-\delta$ (with the required conversions of units). The positions of the two stars at $t_{0}$ and future time $t_{1}=t_{0}+\Delta t$ are given in the Cartesian coordinates by
$\left.\begin{array}[]{ccc}X_{i,0}&=&R_{i}\sin\theta_{i}\cos\phi_{i}\\ Y_{i,0}&=&R_{i}\sin\theta_{i}\sin\phi_{i}\\ Z_{i,0}&=&R_{i}\cos\theta_{i}\end{array}\right\}\text{\hskip 10.00002ptand\hskip 10.00002pt}\left.\begin{array}[]{ccc}X_{i,1}&=&X_{i,0}+V_{X}\Delta t\\ Y_{i,1}&=&Y_{i,0}+V_{Y}\Delta t\\ Z_{i,1}&=&Z_{i,0}+V_{Z}\Delta t\end{array}\right\},$ (B4)
where $i=A,B$. For the two radii, I take $R_{A}=d_{M}$ (the error-weighted mean of the two observed distances) and consider the case that they are the same or $R_{B}=R_{A}+\mathcal{N}(0,2s)$ where $s$ is the observed projected separation between the two stars.
At $t_{1}$, the spherical coordinates of the two stars are given by
$\left.\begin{array}[]{ccc}R_{i,1}&=&\left(X_{i,1}^{2}+Y_{i,1}^{2}+Z_{i,1}^{2}\right)^{1/2}\\ \theta_{i,1}&=&\cos^{-1}\left(Z_{i,1}/R_{i,1}\right)\\ \phi_{i,1}&=&\tan^{-1}\left(Y_{i,1}/X_{i,1}\right)\end{array}\right\}.$ (B5)
The differences in the angular and radial motions of the two stars from $t_{0}$ to $t_{1}$ are then given by
$\left.\begin{array}[]{ccc}v_{\rm{pers},r}&=&\frac{R_{B,1}-R_{B,0}}{\Delta t}-\frac{R_{A,1}-R_{A,0}}{\Delta t}\\ \mu_{\rm{pers},\delta}&=&\frac{-(\theta_{B,1}-\theta_{B,0})}{\Delta t}-\frac{-(\theta_{A,1}-\theta_{A,0})}{\Delta t}\\ \mu^{\star}_{\rm{pers},\alpha}&=&\sin\theta\left(\frac{\phi_{B,1}-\phi_{B,0}}{\Delta t}-\frac{\phi_{A,1}-\phi_{A,0}}{\Delta t}\right)\end{array}\right\},$ (B6)
where $\theta$ refers to the mean polar angle given in Equation (B1). These are the perspective effects of star $B$ with respect to star $A$. To correct for these perspective effects, they are added to the observed RV and PM components of star $A$.
Perspective effects can be a concern only for wide binaries of relatively large separations. Figure 28 shows distributions of perspective effects for 82 wide binaries in a MOND regime ($s/r_{\rm{M}}>0.5$) from the statistical sample (see Table 1). Here perspective effects of PMs are re-expressed as velocities. Because the system velocities (Equation (B2)) are in random directions, the median perspective effects are zero, and only the dispersions are relevant here. For the case $d_{B}=d_{A}$, the standard deviation of all perspective velocities is $37.2$ m s-1, which is comparable to or much smaller than the Gaia DR3 measurement uncertainties of the velocities in the R.A., decl., and radial directions. The perspective velocities normalized by $v$ (which is the magnitude of the observed physical velocity) have a standard deviation of $0.13$. This does not necessarily mean that $v$ is systematically boosted by an appreciable amount because perspective effects can occur in the positive or negative direction for each velocity component. A proper estimate of any systematic bias in the inferred gravity that perspective effects may cause requires individual Bayesian modeling with the PMs and RVs corrected for the perspective effects. Comparison of the second row in Figure 26 with the first row in Figure 22 shows that the effect is very minor (with $\Delta\Gamma=-0.006$ in the MOND regime).
For the case $d_{B}=d_{A}+\mathcal{N}(0,2s)$, the perspective effects have a standard deviation of $\approx 50$ m s-1, which is still comparable to or much smaller than the Gaia DR3 measurement uncertainties of the velocities in the R.A., decl., and radial directions. Moreover, because perspective effects can increase or decrease depending on the distance deviates in individual binaries, the large scatter does not necessarily lead to more reduction of gravitational anomaly than in the case $d_{B}=d_{A}$. To see the effects of the distance deviate $\mathcal{N}(0,2s)$ on $\Gamma$, I have carried out Bayesian modeling of a statistical sample generated by correcting for the perspective effects in PMs and RVs with $d_{B}=d_{A}+\mathcal{N}(0,2s)$. Perhaps ironically, the magnitude of $\Delta\Gamma$ in the MOND regime can be less than in the case $d_{B}=d_{A}$. When the distance difference between the two stars is not measurable, it is perhaps most reliable to assume that it is zero as done in the main body of this paper.
Figure 28: Perspective effects are shown for three relative velocity components due to PMs in R.A. and decl. directions and RVs of stars in wide binaries in the MOND regime with $s/r_{\rm{M}}>0.5$, where $s$ is the sky-projected separation between the two stars and $r_{\rm{M}}$ is the MOND radius. The results shown in the left column are obtained assuming that the two stars are at the same distance from the Sun while the right column shows the results assuming that two distances differ by normal deviates with a standard deviation of $2s$. The upper row shows perspective effects in physical units of m s-1 while the lower row shows the values normalized by physical 3D velocity magnitudes ($v$). Each panel indicates the standard deviation of $N_{\rm{binary}}$ perspective effects.